# The revised boomerang connectivity tables and their connection to the difference distribution table

### v2

@@ 8. DBCT for the Gold function
Let $F_{1}(X)=X^{2^{s}+1}$, over $\mathbb{F}_{2^{n}}$ where $\gcd(s,n)=t$ and $m=\frac{n}{t}$ is odd. Then,
$\operatorname{DBCT}_{F_{1}}(a,d)=\begin{cases}2^{2t}\left((2^{t}-2)+|N(a,d)|\right)&\mbox{~{}if~{}}{\rm Tr}_{s}^{sm}\left(\frac{d}{(a^{2^{s}+1})^{2^{s}+1}}\right)=u^{4},u\in\mathbb{F}_{2^{t}}^{*}\setminus\{1\}\\ 2^{2t}\cdot|N(a,d)|&\mbox{otherwise,}\end{cases}$
@@
First, assume that $a=0$ or $d=0$, then using definition of DBCT, we can write $\operatorname{DBCT}_{F_{1}}(a,d)=2^{2n}$. For $a,d\in\mathbb{F}_{2^{n}}^{*}$, we can write
@@
$\operatorname{LBCT}_{F_{1}}(b,c,d)$ =$=\begin{cases}2^{t}&\mbox{ if }{\rm Tr}_{s}^{sm}\left(\frac{d}{c^{2^{s}+1}}\right)={\rm Tr}_{1}^{m}(1)\mbox{ and }b\in c\mathbb{F}_{2^{t}}^{*}\mbox{ for }b,c,d\in\mathbb{F}_{2^{n}}^{*}\\ \operatorname{DDT}_{F_{1}}(c,d)&\mbox{ if }b=0\\ 0&\mbox{otherwise,}\end{cases}$
$\operatorname{UBCT}_{F_{1}}(a,b,c)$ =$=\begin{cases}2^{t}&\mbox{ if }{\rm Tr}_{s}^{sm}\left(\frac{b}{a^{2^{s}+1}}\right)={\rm Tr}_{1}^{m}(1)\mbox{ and~{}for~{}}a,b,c\in\mathbb{F}_{2^{n}}^{*},\\ &\qquad c\in\left\{b,(u+u^{2})a^{2^{s}+1}+ub,u\in\mathbb{F}_{2^{t}}^{*}\setminus\{1\}\right\}\\ \operatorname{DDT}_{F_{1}}(a,b)&\mbox{ if }c=0\\ 0&\mbox{otherwise.}\end{cases}$
@@
Let us now have $b=c$, then
$\sum_{b=c}\operatorname{UBCT}_{F_{1}}(a,b,c)\operatorname{LBCT}_{F_{1}}(b,c,d)=$ $\sum_{b}\operatorname{DDT}_{F_{1}}(a,b)\operatorname{DDT}_{F_{1}}(b,d),$
when $\sum_{i=0}^{m-1}\frac{b^{2^{si}}}{a^{2^{si}(2^{s}+1)}}=\sum_{i=0}^{m-1}\frac{d^{2^{si}}}{b^{2^{si}(2^{s}+1)}}=m$. If any of these sums is different from $m$, then the product of the DDT entries is 0. Hence, we need to compute $|N(a,d)|$, that is,
@@
To illustrate Theorem 8.1, in Table 5 we give explicit values of the DBCT entries of Gold function over $\mathbb{F}_{2^{n}}^{*}=\langle g\rangle$, where $g$ is a primitive element of $\mathbb{F}_{2^{n}}$.
@@ Acknowledgements
The authors would like to thank the editor, Prof. Daniele Bartoli, for the prompt handling of our paper, and they extend their appreciation to the very professional referees, who spotted errors/typos (now fixed), provided beneficial and constructive comments to improve our paper. In particular, one of the referees challenged us to describe the CCZ/EA/A-equivalence of the considered concepts, which is now part of Section 4.

### v3

@@ 8. DBCT for the Gold function
Let $F_{1}(X)=X^{2^{s}+1}$ over $\mathbb{F}_{2^{n}}$ where $\gcd(s,n)=t$ and $m=\frac{n}{t}$ is odd. Then,
$\operatorname{DBCT}_{F_{1}}(a,d)=\begin{cases}2^{2n}&\mbox{~{}if~{}}ad=0,\\ 2^{2t}\left((2^{t}-2)+|N(a,d)|\right)&\mbox{~{}if~{}}ad\neq 0,{\rm Tr}_{s}^{sm}\left(\frac{d}{(a^{2^{s}+1})^{2^{s}+1}}\right)=u^{4},u\in\mathbb{F}_{2^{t}}^{*}\setminus\{1\},\\ 2^{2t}\cdot|N(a,d)|&\mbox{otherwise,}\end{cases}$
@@
First, assume that $a=0$ or $d=0$, then using definition of DBCT, we can write $\operatorname{DBCT}_{F_{1}}(a,d)=2^{2n}$. Here $F_{1}$ is a permutation as $m$ is given to be odd. Thus, for $a,d\in\mathbb{F}_{2^{n}}^{*}$, we can write
@@
$\operatorname{LBCT}_{F_{1}}(b,c,d)$ =$=\begin{cases}2^{t}&\mbox{if }{\rm Tr}_{s}^{sm}\left(\frac{d}{c^{2^{s}+1}}\right)={\rm Tr}_{1}^{m}(1)\mbox{ and }b\in c\mathbb{F}_{2^{t}}^{*}\mbox{ for }b,c,d\in\mathbb{F}_{2^{n}}^{*},\\ \operatorname{DDT}_{F_{1}}(c,d)&\mbox{if }b=0,\\ 0&\mbox{otherwise,}\end{cases}$
$\operatorname{UBCT}_{F_{1}}(a,b,c)$ =$=\begin{cases}2^{t}&\mbox{if }{\rm Tr}_{s}^{sm}\left(\frac{b}{a^{2^{s}+1}}\right)={\rm Tr}_{1}^{m}(1)\mbox{ and~{}for~{}}a,b,c\in\mathbb{F}_{2^{n}}^{*},\\ &\qquad c\in\left\{b,(u+u^{2})a^{2^{s}+1}+ub,u\in\mathbb{F}_{2^{t}}^{*}\setminus\{1\}\right\},\\ \operatorname{DDT}_{F_{1}}(a,b)&\mbox{if }c=0,\\ 0&\mbox{otherwise.}\end{cases}$
@@
If $b=c$, then $\sum_{b=c}\operatorname{UBCT}_{F_{1}}(a,b,c)\operatorname{LBCT}_{F_{1}}(b,c,d)=\sum_{b}\operatorname{DDT}_{F_{1}}(a,b)\operatorname{DDT}_{F_{1}}(b,d),$ when $\sum_{i=0}^{m-1}\frac{b^{2^{si}}}{a^{2^{si}(2^{s}+1)}}=\sum_{i=0}^{m-1}\frac{d^{2^{si}}}{b^{2^{si}(2^{s}+1)}}=m$. If any of these sums is different from $m$, then the product of the DDT entries is 0. Hence, we need to compute $|N(a,d)|$, that is,
@@
To illustrate Theorem 8.1, in Table 5 we give explicit values of the DBCT entries of the Gold function over $\mathbb{F}_{2^{n}}^{*}=\langle g\rangle$, where $g$ is a primitive element of $\mathbb{F}_{2^{n}}$.
@@ Acknowledgements
The authors would like to thank the editor, Prof. Daniele Bartoli, for the prompt handling of our paper, and they extend their appreciation to the very professional referees, who spotted errors/typos (now fixed), and provided beneficial and constructive comments to improve our paper. In particular, the referees challenged us to describe the CCZ/EA/A-equivalence of the considered concepts, which is now part of Section 4, or provide examples, whenever possible.